# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

### v1

@@ 3.2.1 Preconditioning
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, and $\mathcal{F}$ is the transfer factor. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. Here there are two important temperatures: chamber wall temperature $T_{\textrm{c}}$ and cold gas temperature $T_{\textrm{g}}$. The top surface (chamber wall) and local gas in that area are not cooled, so we assume the gas temperature is equal to the wall temperature $T_{\textrm{c}}$, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection. Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and thus $h_{\textrm{s3}}$ takes into account of the convection part only.
@@
As mentioned before, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by Ref. [6] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6).
3.2.2 Vacuum-induced surface freezing
@@
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until VISF completes at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure of a system to evaporate the liquid solution, abruptly decreasing the temperature and promoting nucleation. A model for VISF can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref. [31]. First, consider the mass transfer part. The evaporation rate of water at the surface for nonvolatile solute is
$\frac{dm_{\textrm{w}}}{dt}=h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (7)
@@
The initial condition for (7) is simply the final temperature of the preconditioning step.
@@
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}(t)+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (11)
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization. At low temperature, $\Delta H_{\textrm{vap}}$ can be approximated by
@@
In (11), $m_{\textrm{w}}(t)$ and $Q_{\textrm{s3}}(t)$ are time-dependent because the amount of water (and hence the side surface area $A_{r}$) changes with time. The initial condition for (11) is the final mass of water at the end of the preconditioning step, which is $m_{\textrm{w,0}}$.
Note that, if the VISF process is done properly, the amount of water vaporized is usually very small and so can be safely neglected.
3.2.3 Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our strategy to treat the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref. [11]. In this work, we use the term nucleation to denote the first nucleation where the temperature of a subcooled liquid almost instantaneously increases to the equilibrium/freezing point, whereas the term solidification is used to denote the phase transition from liquid to solid after that first nucleation.
Consider the controlled nucleation case first. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
@@ 3.2.3 Nucleation
where $T_{\textrm{f,l}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (13) and (14) simultaneously, both of which become the initials conditions for the next (solidification) step.
@@
For the uncontrolled nucleation case (no VISF), the entire calculation is Section 3.2.2 is skipped. By nature, ice nucleation is stochastic and interpreted as a Poisson process. The rate constant of the Poison process $\lambda$ is [32]
@@
The probability defined by (22) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. At this point, the exact same procedure described for the controlled nucleation case, i.e.,(13)–(19), is followed, in which $t_{\textrm{f2}}$ is replaced with $t_{\textrm{f1}}$.
3.2.4 Solidification
After the first nucleation completes, the solidification step starts at $t_{\textrm{f3}}$. During this step, the product temperature is equal to the freezing point of the solution, which follows the freezing-point depression relation
$T(t)=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}(t)}\right),$ (23)
The energy balance during the freezing step is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}}+m_{\textrm{i}}(t)C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}(t)+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (24)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. During the solidification step, the amount of water and ice changes with time, which affects the volume and surface area of the product, and so $Q_{\textrm{s2}}(t)$ is time-dependent. The relation between the amount of water and ice is
$m_{\textrm{w}}(t)=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (25)
@@ 3.2.4 Solidification
3.2.5 Cooling
@@
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, taking the final conditions of the solidification step. The initial condition for (27) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (27) without any significant error.
By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature and amount of substances (e.g., water and ice) during the freezing
@@
In conventional lyophilization where vials are placed on the heating shelf, there are three heat transfer mechanisms associated with the bottom surface, namely conduction, convection, and radiation, whereas only radiation dominates at the side and top surfaces [17]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated following the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
The governing equations for the primary drying step consists of (1) the energy balance in the frozen region and (2) the mass balance at the sublimation front. By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region is
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S(t)<z<H,$ (28)
where $T(z,t)$ is the temperature, $S(t)$ is the sublimation front/interface position, $k$ is the thermal conductivity, $\rho$ is the density, $C_{p}$ is the heat capacity, $V$ is the volume, $H$ is the height of the product, and the subscript ‘f’ denotes the frozen region. We refer to Appendix A for the calculations of some relevant parameters. The radiative heat transfer from the sidewall $Q_{\textrm{rad}}$ is
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (29)
where $\mathcal{F}$ is the transfer factor, $\sigma$ is the Stefan-Boltzmann constant, $T_{\textrm{c}}$ is the chamber wall temperature, and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of (29). The mass balance of water at the sublimation front gives
@@
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation can be is described by [22]
@@
The variation of the resistance normally follows the empirical expression [17, 21, 22]
@@
where $R_{\textrm{p}}$ is the mass transfer resistance (aka cake resistance) and $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, $R_{\textrm{p2}}$ are the constants estimated from data.
The partial differential equation (PDE) (28) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H,$ (34)
where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient which combines the effects of thermal radiation and natural convection. At the top surface, the energy balance associated with sublimation and thermal radiation from the above shelf/wall is
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=S(t),$ (35)
where $\Delta H_{\textrm{sub}}$ is the heat of sublimation and $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial.
@@
If the model is simulated following the freezing step, then set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
The primary drying model is simulated until the interface position is equal to the height of the product $S(t)=H$, indicating that there is no frozen material left, which marks the end of the primary drying step $t_{\textrm{d1}}$. In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying stage. However, this period is trivial to model; only a simple heat equation is required, and so is not described it here.
@@
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) as done for the primary drying model (Fig. 3C). The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated following the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
The governing equations for the secondary drying step consists of (1) the energy balance in the dried region and (2) the desorption kinetics. The energy balance of the dried product is
@@
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the bound water concentration (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the enthalpy of desorption, $Q_{\textrm{rad}}$ is as defined in (29), and the subscript ‘e’ denotes the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
@@
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{w}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [14, 15, 21]
@@
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency) and $E_{\textrm{a}}$ is the activation energy. The above desorption kinetics (39) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [14, 38, 18, 23, 21]. Further simplification that is very common in the literature can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in Ref. [15] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [21].
@@
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H.$ (41)
Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=0.$ (42)
@@
If the model is simulated following the primary drying step, then set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step $t_{\textrm{d2}}$.
@@
Convection and radiation are two key heat transfer mechanisms in the suspended-vial configuration as discussed before. During the freezing step, there could be a combination of natural/forced convection and radiation, depending on how the cryogenic gas is fed to the system. In primary and secondary drying, only natural convection and radiation are important.
@@
Convection can be simply modeled using Newton’s law of cooling. A heat transfer coefficient for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref. [21].

### v2

@@ Introduction
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, $\mathcal{F}$ is the transfer factor (see the definition in Section 3.5), and the subscripts ‘u’, ‘b’, ‘c’, and ‘g’ denote the upper surface of the chamber, bottom shelf, chamber walls, and cold gas, respectively. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. In lyophilization, the bottom shelf and gas temperatures ($T_{\textrm{b}}$ and $T_{\textrm{g}}$) typically vary with time and are specified by the freezing/drying protocol. The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [39]. In this article, we assume that both temperatures are approximately constant as reported by Refs. [40, 19]. In addition, $T_{\textrm{u}}\approx T_{\textrm{c}}$ due to no additional heat source/sink at these locations. Nevertheless, our model is not restricted to these assumptions; i.e., temperature-dependent data can be fed to the model (if available).
Next, (4)–(6) are simplified further to match the experimental setup in Ref. [7], whose data are used for our model validation. In Ref. [7], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection. In this setup, the upper surface and local gas in that area are not cooled, so the gas temperature is assumed to be equal to the surface temperature $T_{\textrm{u}}$. Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of Ref. [31] and Section 3.5). Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. As a result, $Q_{\textrm{s1}}$ and $Q_{\textrm{s2}}$ are
$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{u}}-T),$ (7)
$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T).$ (8)
At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for (6).
@@
Note that the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by Ref. [7] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6) even further.
Vacuum-induced surface freezing
@@
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until its completion at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure to evaporate a small amount of liquid solution from the product, abruptly decreasing the temperature and promoting nucleation. A model for VISF is not readily available in the literature but can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref. [31]. First, consider the mass transfer part. The evaporation rate of water at the surface for a nonvolatile solute is
$\frac{dm_{\textrm{w}}}{dt}=-h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (9)
@@
Note that, if the VISF process is done properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
@@
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (13)
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [41]
@@
In (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
The initial conditions for (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref. [12]. In this work, we use the term nucleation to denote the first nucleation where the temperature of supercooled liquid almost instantaneously increases to the equilibrium/freezing point, which is caused by the heat released from the fraction of liquid being frozen/solidified. On the other hand, we use the term solidification to denote the phase transition from liquid to solid (ice formation) after that first nucleation.
First, consider the case of controlled nucleation with VISF. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
@@
where $T_{\textrm{f,w}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (15) and (16) simultaneously.
@@
For the case of stochastic nucleation, the entire calculation described for VISF in Section 3.2.2 is skipped. By nature, ice nucleation is stochastic and can be interpreted as a Poisson process. The rate constant of the Poison process $\lambda$ is
@@
The probability defined by (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. Subsequently, follow the exact same procedure described by (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
Solidification
After the first nucleation completes, the solidification (aka ice formation) step starts at $t_{\textrm{f3}}$. For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [12]; we also discuss this in the context of the Biot number in Appendix A. The lumped capacity model is sufficiently accurate and highly computationally efficient, but it does not represent the correct physics of ice formation. By treating the entire solid/liquid as a lumped object during solidification, it implies that homogeneous nucleation is assumed. Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [42]. Consequently, the solidification process generally starts from the vial surfaces, i.e., the outer and bottom surfaces in this case. To take this fact into account, a rigorous 2D heat transfer model could be considered [13]. In such cases, the physics is incorporated more accurately, but the major drawback of 2D modeling is its high computational cost and numerical complexity.
Our strategy here is to combine the advantages of the lumped capacity and 2D models together, resulting in a model that is computationally efficient, provides accurate prediction, and represents the correct physics; we denote it as a hybrid lumped capacity model. In this case, ice formation is assumed to initiate from the side and bottom surfaces (similar to a 2D model), in which the ice layer is treated as an additional heat transfer resistance between the cold gas and unfrozen water. Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [31]. Physically, this is equivalent to assuming that all the heat input is used for phase transition only. The unfrozen water and solute are treated as a lumped object as its temperature follows the equilibrium temperature. A schematic diagram showing our model for the solidification step is shown in Fig. 4
Figure 4: Schematic diagram showing the mechanistic modeling of the solidification step. For simplification, it is assumed that the liquid part retains a cylindrical shape with the same aspect ratio as the initial solution before nucleation starts.
The energy balance during the solidification step is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (25)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively. During the solidification step, the amount of water $m_{\textrm{w}}(t)$ and ice $m_{\textrm{i}}(t)$ varies with time, which also affects the volume and surface area of the product. The relation between $m_{\textrm{w}}$ and $m_{\textrm{i}}$ is
$m_{\textrm{w}}=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (26)
The temperature of the liquid phase and the solid-liquid interface follows the freezing-point depression relation
$T=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}}\right).$ (27)
The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in (8) and (6) can be replaced by the overall heat transfer coefficients
$U_{\textrm{s2}}=\frac{1}{1/h_{\textrm{s2}}+l/k_{\textrm{i}}},$ (28)
$U_{\textrm{s3}}=\frac{1}{1/h_{\textrm{s3}}+\frac{r_{\textrm{o}}\ln(r_{\textrm{o}}/r)}{k_{\textrm{i}}}}.$ (29)
where $r_{\textrm{o}}$, $r$ and $l$ are as defined in Fig. 4 and $k_{\textrm{i}}$ is the thermal conductivity of ice. For the radiation part, the linearization technique described in Section 3.2.1 can be applied, and so the equation is written in the form of Newton’s law of cooling. Note that the overall heat transfer coefficients in (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$
@@
Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in Ref [12]) while describing the physics in a more accurate way as in the 2D model (e.g., in Ref [13]). Final simulations results from these models are almost identical for lyophilization of unit doses, and so there is no concern in terms of model accuracy.
Cooling
@@
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, following the final conditions of the solidification step. The initial condition for (31) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (31) without any significant error.
By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature, phase transition, and amount of ice/water during the freezing step can be predicted.
@@
In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [18]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated consecutively after the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
The governing equations for the primary drying step consist of the (1) energy balance in the frozen region and (2) mass balance at the sublimation front. By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region can be described by the partial differential equation (PDE)
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S<z<H,$ (32)
where $T(z,t)$ is the temperature, $S(t)$ is the sublimation front/interface position, $k$ is the thermal conductivity, $\rho$ is the density, $C_{p}$ is the heat capacity, $V$ is the volume, $H$ is the height of the product, and the subscript ‘f’ denotes the frozen region. We refer to Appendix B for the calculations of some relevant parameters. The radiative heat transfer from the sidewall $Q_{\textrm{rad}}$ is
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}_{\textrm{s3}}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (33)
where $T_{\textrm{c}}$ is the chamber wall temperature and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of (33). The mass balance of water at the sublimation front gives
@@
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation is described by [22]
@@
The variation of the mass transfer resistance can be approximated by the empirical expression [18, 4, 22]
@@
where $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, and $R_{\textrm{p2}}$ are the constants to be estimated from data. Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
The PDE (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient that combines the effects of thermal radiation and natural convection. At the top surface, the energy balance associated with sublimation and thermal radiation from the upper surface of the chamber is
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=S,$ (39)
where $\Delta H_{\textrm{sub}}$ is the heat of sublimation.
@@
For consecutive simulation with the freezing step, set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
The primary drying model is simulated until the interface position is equal to the height of the product, i.e., $S=H$, indicating that there is no frozen material left, which marks the end of the primary drying step at $t_{\textrm{d1}}$. In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying step. However, the model contains only a simple heat equation, and so it is not detailed here.
@@
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3C), which is consistent with the primary drying model. The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated consecutively after the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
The governing equations for the secondary drying step comprise the (1) energy balance in the dried region and (2) desorption kinetics. The energy balance of the dried product is
@@
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in (33), and the other parameters are as defined in (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
@@
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [15, 16, 4]
@@
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency), $E_{\textrm{a}}$ is the activation energy, and $R$ is the gas constant. The above desorption kinetics (43) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [15, 43, 19, 23, 4]. Further simplification that is relatively common can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in Ref. [16] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [4].
@@
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying (38). Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=0.$ (46)
@@
For consecutive simulation with the primary drying step, set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step at $t_{\textrm{d2}}$.
@@
Convection and radiation are the two important modes of heat transfer in suspended vials as discussed extensively in the previous sections. During the freezing step, there could be a combination of natural/forced convection and radiation, depending on how the cryogenic gas is fed to the system. In primary and secondary drying, only natural convection and radiation are important.
@@
Convection can be simply modeled using Newton’s law of cooling. Heat transfer coefficients for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref. [4].